I was trying the tutorial and made it to the point where it says just type in the equation to solve it and it just wont solve this is not for school just trying to learn how much solid wrap and printed wrap was sold what i got was
4.00x + 6.00y = 2340
x + y = 480

Found 2 solutions by Alan3354, ankor@dixie-net.com:
Answer by Alan3354(69443)

4x + 6y = 2340 --> 2x + 3y = 1170
x + y = 480 --> y = 480-x
Sub for y into eqn 1
2x + 3(480-x) = 1170
-x = 1440 = 1170
x = 270
y = 210

Answer by ankor@dixie-net.com(22740)

You can use substitution
4.00x + 6.00y = 2340
:
x + y = 480
y = (480-x)
:
Substitute for y in the 1st equation and find x
4x +6(480-x) = 2340
4x + 2880 - 6x = 2340
4x - 6x = 2340 - 2880
-2x = -540
x = 270
:
y = 480 - 270
y = 210
;
;
Check solutions in the 1st equation
4(270) + 6(210) =
1080 + 1260 = 2340